# A Benchmark for Optimal Multi-Modal Multi-Robot Multi-Goal Path Planning with Given Robot Assignment

### v1

@@ VII Conclusion
We presented a formalization of the multi-modal multi-robot multi-goal motion planning problem, and a benchmark containing diverse problems reaching from simple environments with short goal sequences, to relatively long horizon problems requiring coordination of multiple robots. We introduced two planners that are probabilistically complete, with both of them being asymptotically optimal, based on an RRT planner, and a PRM planner respectively.
The planners that we propose are not expected to scale to very long horizon problems or problems with a large amount of robots. However, we believe that considering the problem in composite space helps understand the implication that, e.g., priorization in planning has in the continuous space, and through that better understanding, help develop better planners.
There are some clear improvements possible, such as such exploiting the multi-robot structure of the planning problems by adapting, e.g., dRRT to our problems. Further, the nature of the mode families [39], or the multi-query-like structure of the problem [40] could be used. We also believe that receding horizon-planners are an interesting avenue of research in order to transition this purely offline planning approach to real world execution.
Finally, the formulation we propose supports extension of the transition logic to multi-robot task and motion planning by changing how the next mode, respectively the mode sequence is generated.
@@ Acknowledgments
The authors thank Miguel Zamora and Yijiang Huang for insightful discussions and comments on both text and illustrations.

### v2

@@ VII Conclusion
We presented a formalization of the multi-modal, multi-robot, multi-goal motion planning problem, and open-source the benchmark we developed, containing diverse problems reaching from simple environments with short goal sequences, to relatively long horizon problems requiring coordination of multiple robots. We adapt multiple standard sampling-based planners that are probabilistically complete and asymptotically optimal to this type of problem. We also benchmarked a prioritized planner against these optimal planners and show that the solutions found by such a suboptimal planner can be competitive in some settings.
There are some clear improvements possible, such as exploiting the multi-robot structure of the planning problems by adapting, e.g., dRRT to our problems. Further, the nature of the mode families [43], or the multi-query-like structure of the problem [44] could be used. We also believe that receding horizon-planners are an interesting avenue of research in order to transition this purely offline planning approach to real-world execution.
There are many extensions to the chosen problem formulation: The formulation and planners we propose supports extension of the transition logic to full multi-robot task and motion planning by changing the task-assignment-oracle. In the future, we also want to support constrained multi-robot planning, or kinodynamic planning.
@@ Acknowledgments
The authors thank Miguel Zamora for insightful discussions and comments on both text and illustrations.